a carpenter can construct bookcases at a cost of 40 each. if the carpenter sells the bookcases for x dollars each, it is estimated that 300-2x bookcases will be sold per month. determine the selling price of each bookcase that will give the carpenter the greatest monthly profit.
Answer by stanbon(75887)

a carpenter can construct bookcases at a cost of 40 each. if the carpenter sells the bookcases for x dollars each, it is estimated that 300-2x bookcases will be sold per month. determine the selling price of each bookcase that will give the carpenter the greatest monthly profit.
------
Cost to produce 300-2x bookcases = 40(300-2x) = 12000-80x dollars
--
Revenus when 300-2x bookcases are sold : (300-2x)x = 300x - 2x^2
-----
Profit = (revenue-cost) = -2x^2 + 380x - 12000
----
Maximum profit occurs when x = -b/(2a) = -380/(2*-2) = 95
---------------
Cheers,
Stan H.
===============